if the marginal revenue on each MRI is $2500, and the marginal cost for each is $500, what would be the profit or loss on 52 MRI's performed at your clinic last month? a. 118,000 b. 156,000 c. 104,000 d. loss of $26,000
Answer #1

The answer is 104000

Marginal revenue on each MRI = 2500

Marginal Cost on each MRI = 500

so total Revenue at 52 MRI = marginal revenue in each MRI * No of MRI

= 2500 * 52 = 130000

Total cost at 52 MRIs = Marginal Cost in each MRI * No of MRI

= 500 * 52 = 26000

Profit = TR - TC = 130000 - 26000 = 104000